WHICH LIGHT BULB IS THE BRIGHTEST? CIRCLE THE LIGHT BULB THAT IS THE BRIGHTEST 40W 100W 25W Ri P; PI 1ー 120V I. Calculate the resistance for euch bulb at 120V for the stated power and enter in boxes above 2. Calculate Rsg for the circuit. 3. Calculate the current in the circuit and enter in the box above 4. Calculate the voltage drop across each bulb and enter in boxes above. 5. Calculate the power consumed by euch bulb and enter in boxes above.
Answer #1

Please see There are two scanned sheets attached.

First of all Lets discuss Which bulb is brightest:

Since all bulbs are connected in series therefore current flowing through each bulb is same. Since Resistances of each bulb are different therefore the brightness is different because brightness depends on heat produced or power since heat produced is proportional to resistance hence greater the resistance greater is the brightness.

Resistance of first bulb is greater hence the first bulb of power 25 Watt will be brightest.

Now i am entering the answers in the boxes as asked. The explanations for those answers are attached on a scanned sheet.

  righte to W2 576 Ohm esistanceef 40W bulb 2. Resistanve lao 2. ー JD Res for the ciruu : eq= 576十360+144 = 1080 Ohm eqiven1080 Ohm, 12o current in the circuit L: Re: Togo = 0-111 Am Cummenh in the circuit is o Voltage dro acress each, buub Am pese nsesies combinatiom hene ower cmsumedb each bulb 1s different becaus e-their. Yes Istqnceg are difterent Power consumed b fi따 bulbこ@tu)x576 = 7-10% nhtt atiom hente owe Pwer can sumal ky Second, (0.111) 360 = 4.443.5 watt =(0.111) fewer cønsynd by thi져 44-1.777 nett